OpenStudy (cometailcane):

Calculate how many moles of MnO4-(aq) were present in the volume of FA3 calculated in b. b= 35.3

OpenStudy (cuanchi):

It is missing some information. What is FA3? b=35.3 what are the units of b=35.3?

OpenStudy (cometailcane):

FA3 is 1.00 mol dm-3 sulfuric acid and b is 35.3 cm3 of FA3

OpenStudy (anonymous):

Hey. Long time no see. Alright then, for solution, we use the formul n=cv/1000, where v is volume in cm^3, c is concentration and n is moles. hence n=1*35.5/1000= 0.0355mol.